OpenStudy (anonymous):

Is the answer to 3p+8=-2(p-22) equals 36? I'm confused

OpenStudy (goose.):

No, that's not actually.

OpenStudy (goose.):

5p=36

OpenStudy (goose.):

Divide both sides by 5, to get your simplified answer,

Mehek (mehek14):

\(3p+8=-2p+44\\5p=36\\36/5=p\)

OpenStudy (anonymous):

Thank you all very much.